# Recovery of rational functions via Hankel pencil method and sensitivities of the poles

### v2

@@ 5 Numerical experiments
In this section, we present a few numerical experiments. The algorithms are implemented in Matlab and use IEEE standard floating point arithmetic with double precision. By $\tilde{z}_{j}$ and $\tilde{\gamma}_{j}$ we denote reconstructed poles and coefficients, respectively, and by the following formulas
@@
In Examples 5.1-5.3, we demonstrate recovery results and analyze sensitivities of poles with the method developed in Section 4. We employ $2N$ (noisy) Fourier coefficients $\hat{r}_{-k}(1+\varepsilon_{k})$, $k=1,\ldots,2N$, to reconstruct poles $z_{j}^{(-)}$, and $2N$ noisy Fourier coefficients $\hat{r}_{k}(1+\varepsilon_{k})$, $k=1,\ldots,2N$, to reconstruct poles $z_{j}^{(+)}$. By $(\varepsilon_{k})_{k=1}^{2N}$ we denote the random noise that is generated in Matlab by $\sigma$*randn(2N), i.e., we take multiplicative noise drawn from the standard normal distribution. We take $L=N$ and consider different values of $N$ and $\sigma$. To compute sensitivities $\rho_{j}^{(-)}$, $\rho_{j}^{(+)}$ and ,$\eta_{j}^{(-)}$, $\eta_{j}^{(+)}$, we use formulas (4.14), (4.20) and (4.27), (4.28), respectively. Upper bounds for sensitivities $\rho_{j}^{(-)}$ and $\rho_{j}^{(+)}$ are computed via (4.15) and (4.21), respectively.
@@
min $1.78$e–$04$ $1.17$e–$04$ min $4.03$e–$02$ $2.77$e–$02$
max $2.11$e–$03$ $2.46$e–$03$ max $3.04$e–$01$ $1.44$e–$01$
average $1.44$e–$03$ $8.48$e–$04$ average $1.11$e–$01$ $9.55$e–$02$
Table 3: Reconstruction errors $|z-z_{j}|$ for each pole $z_{j}$ for Example 5.2 for $\sigma=10^{-3}$
@@
min $1.71$e–$05$ $1.13$e–$06$ $1.31$e–$05$ $8.74$e–$06$
max $2.86$e–$04$ $2.7$e–$04$ $1.89$e–$04$ $1.9$e–$04$
average $1.44$e–$04$ $1.19$e–$04$ $1.04$e–$04$ $9.43$e–$05$
Table 4: Reconstruction errors $|z-z_{j}|$ for each poles $z_{j}$ for Example 5.2 for $\sigma=10^{-1}$
@@
min $4.2$e–$03$ $4.25$e–$01$
max $1.73$e–$02$ $1.22$e–$00$
average $9.7$e–$03$ $6.35$e–$01$
@@
4 $-1.1\mathrm{i}$ 0.035 2.208 3.559
@@
Let a rational function $r(z)$ in (1.1) be given by parameters $M=2$ and
$\boldsymbol{z}=(-0.1,-2.1),\quad\boldsymbol{\gamma}=(0.5,0.5)$
(see [6, 23]). When $\sigma=0$, i.e. in case of recovery from exact data, we get the following reconstruction errors with $N=2$,
[…]
For $\sigma=10^{-3}$ and $\sigma=10^{-1}$, we compute 10 iterations and present the results in Table 1. Since we are interested to investigate which pole is more sensitive to perturbations, we compute reconstruction errors for poles $z_{1}^{(-)}=-0.1$ and $z_{1}^{(+)}=-2.1$ separately. Results for the recovery of the coefficient vector are presented in Table 2.
Table 7: Reconstruction errors $|z-z_{j}|$ for each pole $z_{j}$ for Example 5.3 for $N=4$
$\sigma=10^{-7}$ $z_{1}^{(-)}=0.2$ $z_{2}^{(-)}=0.5$ $z_{1}^{(+)}=2$ $z_{2}^{(+)}=50$
min $1.58$e–$08$ $2.9$e–$08$ $6.07$e–$09$ $1.75$e–$03$
max $5.1$e–$07$ $2.49$e–$07$ $2.7$e–$07$ $3.21$e–$01$
average $2.48$e–$07$ $1.3$e–$07$ $1.45$e–$07$ $1.34$e–$01$
$\sigma=10^{-5}$ $z_{1}^{(-)}=0.2$ $z_{2}^{(-)}=0.5$ $z_{1}^{(+)}=2$ $z_{2}^{(+)}=50$
min $4.24$e–$06$ $1.18$e–$06$ $2.57$e–$06$ $3.16$e–$00$
max $6.45$e–$05$ $2.77$e–$05$ $2.92$e–$05$ $9.46$e+$01$
average $2.31$e–$05$ $1.13$e–$05$ $1.5$e–$05$ $2.02$e+$01$
Analyzing numerical results in Table 1, we conclude that for $\sigma=0.1$ our method still reconstructs the locations of the poles quite good. On the other hand, using more (noisy) Fourier coefficients also gives us better recovery results. We notice that a pole $z_{1}^{(-)}=-0.1$ is less sensitive to perturbations than a pole $z_{1}^{(+)}=-2.1$. This numerical observation is also confirmed with sensitivities values. For a pole $z_{1}^{(-)}=-0.1$, we have $\rho_{1}^{(-)}=2\cdot 10^{-1}$ and $\eta_{1}^{(-)}=1.97$, and for a pole $z_{1}^{(+)}=-2.1$, we have $\rho_{1}^{(+)}=9.52\cdot 10^{-1}$ and $\eta_{1}^{(+)}=6.65$. Both unstructured and structured sensitivities are larger for a pole $z_{1}^{(+)}=-2.1$ located outside the unit circle. We compute also vectors of structured perturbations
$(\boldsymbol{\mathcal{S}}^{(1)}_{2})^{(-)}=(2.0\cdot 10^{-1},2.0)^{T}\text{ and }(\boldsymbol{\mathcal{S}}^{(1)}_{2})^{(+)}=(4.2,8.82)^{T},$
which means that a pole $z_{1}^{(-)}=-0.1$ is the most sensitive with respect to the perturbation of the Fourier coefficient $\hat{r}_{-2}$, and a pole $z_{1}^{(+)}=-2.1$ with respect to the perturbation of the Fourier coefficient $\hat{r}_{2}$.
@@
We consider an example when poles are located very close to the unit circle. Let a rational function $r(z)$ of type $(7,8)$ in (1.1), i.e with $M=8$, be given by vectors of poles and coefficients
$\boldsymbol{z}=(0.9,-0.9,0.9\mathrm{i},-0.9\mathrm{i},1.1,-1.1,1.1\mathrm{i},-1.1\mathrm{i}),\quad\boldsymbol{\gamma}=(1,2,3,4,5,6,7,8),$
respectively. First, we consider recovery from the exact data, i.e for $\sigma=0$. For $N=8$, we get the following reconstruction errors
@@
min $7.38$e–$04$ $3.79$e–$02$
max $7.38$e–$03$ $6.95$e–$01$
average $4$e–$03$ $4.11$e–$01$
In case of recovery from noisy data, we compute 10 iterations and present recovery result for each pole $z_{j}$ in Tables 3 and 4. Analyzing numerical results, we conclude that the pole $z_{1}^{(-)}=0.9$ is the most sensitive to perturbations. This conclusion is confirmed with the sensitivities presented in Table 6. Results for the recovery of the coefficient vector are presented in Table 5.
Table 10: $\ell_{p}$-norm and upper bound of $\ell_{p}$-norm of $\boldsymbol{\rho}^{(-)}$ and $\boldsymbol{\rho}^{(+)}$ for Example 5.3 for $p=1,2$
@@
$N=4$ $4.01$e+$01$ $5.725$e+$01$ $2.209$e+$03$ $6.802$e+$03$
Example 5.3.
Finally, let a rational function $r(z)$ of type $(3,4)$ in (1.1), i.e with $M=4$, be given by vectors of poles and coefficients
$\boldsymbol{z}=(0.2,0.5,2,50),\quad\boldsymbol{\gamma}=(1,1,1,1),$
Then we consider recovery from noisy data with $\sigma=10^{-7}$ and $\sigma=10^{-5}$. We compute 10 iterations and present the corresponding results in Table 7. Analyzing numerical results, we can conclude that the pole $z_{2}^{(+)}=50$ is the most sensitive to perturbations. For $\sigma=10^{-5}$ the method does not reconstruct location of this pole correctly. This numerical observation is also predicted by the sensitivities values presented in Table 8. Both unstructured $\rho_{2}^{(+)}=2.204\cdot 10^{3}$ and structured $\eta_{2}^{(+)}=1.577\cdot 10^{4}$ sensitivities are sufficiently larger than sensitivities of other poles. Note, that the locations of the poles $z_{1}^{(-)}=0.2$, $z_{2}^{(-)}=0.5$ and $z_{1}^{(+)}=2$ are reconstructed even for $\sigma=10^{-1}$. We present results regarding recovery of the coefficient vector in Table 9.
We compute also vectors of structured perturbations for poles inside the unit circle
$(\boldsymbol{\mathcal{S}}^{(1)}_{4})^{(-)}=(0.555,4.999,13.333,11.111)^{T},\quad(\boldsymbol{\mathcal{S}}^{(2)}_{4})^{(-)}=(0.222,2.666,9.999,11.111)^{T},$
and outside the unit circle
$(\boldsymbol{\mathcal{S}}^{(1)}_{4})^{(+)}=(3.472\cdot 10^{-3},0.354,9.375,17.36)^{T},$
and
$(\boldsymbol{\mathcal{S}}^{(2)}_{4})^{(+)}=(54.25,2.929\cdot 10^{3},1.106\cdot 10^{4},1.085\cdot 10^{4})^{T}.$
That means that poles $z_{1}^{(-)}=0.2$ and $z_{2}^{(-)}=0.5$ are the most sensitive with respect to perturbations of the Fourier coefficient $\hat{r}_{-3}$, the pole $z_{1}^{(+)}=2$ – the Fourier coefficient $\hat{r}_{4}$, and the pole $z_{2}^{(+)}=50$ – the Fourier coefficient $\hat{r}_{3}$.
In Subsection 4.2, we derived the upper bounds (4.224.24) for the $\ell_{2}$- and the $\ell_{1}$-norms of vectors $\boldsymbol{\rho}^{(-)}$ and $\boldsymbol{\rho}^{(+)}$ and we want to understand how sharp these estimates are. In order to do it, we compute the exact values for the corresponding norms and the upper bounds obtained in (4.224.24) and present the corresponding results in Table 10. Comparing these results we can conclude that the estimates (4.224.24) are quite sharp.
Table 11: Reconstruction results for $r$ from Example 5.1
@@
min $4.34$e–$03$ $1.06$e–$03$ min $2.37$e–$02$ $2.27$e–$02$
max $1.23$e–$02$ $4.24$e–$03$ max $2.05$e+$00$ $7.56$e–$02$
average $7.64$e–$03$ $2.71$e–$03$ average $4.52$e–$01$ $4.55$e–$02$
Table 12: Reconstruction results for $r$ from Example 5.2
@@
min $2.2$e–$03$ $2.42$e–$03$ min $1.54$e–$02$ $9.73$e–$03$
max $1.37$e–$02$ $4.86$e–$02$ max $6.69$e–$02$ $3.54$e–$02$
average $6.93$e–$03$ $1.06$e–$02$ average $4.42$e–$02$ $5.75$e–$02$
In Example 5.4 we present the comparison results of the performance of our approach with the AAA method for rational approximation. We employ $4N$ (noisy) function values $r(t_{k})(1+\varepsilon_{k})$ with $t_{k}=\mathrm{e}^{\frac{2\pi\mathrm{i}k}{4N}}$ for $k=0,\ldots,4N-1$, as input information for both methods. By $(\varepsilon_{k})_{k=0}^{4N-1}$ as above we denote the random noise that is generated in Matlab by $\sigma$*randn(4N), i.e., we take multiplicative noise drawn from the standard normal distribution. The corresponding modification of our method is discussed in Remark 4.1. We referee the reader to [14] regarding description of the AAA method. Note only that AAA takes the form of iteration with respect to a type of a rational approximant that is represented at each step in the barycentric form. In case of reconstruction of rational function of type $(M-1,M)$ using exact data, AAA terminates after $M+1$ iteration steps and returns $r$ as
@@
where points $t_{k_{s}}$ are chosen by some greedy algorithm and weights $w_{s}$ are computed by solving a linear least-squares problem over the subset of sample points that have not been selected as support points, $\{t_{0},\ldots,t_{4N-1}\}\setminus\{t_{k_{1}},\ldots,t_{k_{M+1}}\}$ (see also [8, 9, 10] for details). The AAA method has the overall computational cost of $\mathcal{O}(NM^{3})$.
Table 13: Reconstruction results for $r$ from Example 5.3
@@
We consider again recovery of rational functions $r$ from Examples 5.1-5.3. We reconstruct $r$ from $4N$ samples $r(t_{k})(1+\varepsilon_{k})$, $k=0,\ldots,4N-1$ by using our method based on Hankel pencil and AAA. In Tables 11- 13 we present the results for $N=15$, $L=20$ and different values of $\sigma$. We see that for the rational function from Example 5.2 both methods work very well.
Analysing results of Tables 11 and 13, we conclude that for $r$ from Examples 5.1 and 5.3 we get a slightly better recovery from noisy input data with our method. The reason for this is that the method based on the Hankel pencil works better for the recovery of poles with certain locations. For Example 5.3 the poles $z_{1}^{(-)}=0.2$, $z_{2}^{(-)}=0.5$ and $z_{1}^{(+)}=2$ are reconstructed equally good by both methods, while the Hankel pencil based approach shows better performance for the recovery of the pole $z_{2}^{(+)}=50$ than the AAA method (see Tables 14).
Table 14: Reconstruction errors $|z-z_{j}|$ for each pole $z_{j}$ for Example 5.3 with Hankel pencil and the AAA methods

### v3

@@ 5 Numerical experiments
In this section, we present several numerical experiments. The algorithms are implemented in Matlab and use IEEE standard floating point arithmetic with double precision. By $\tilde{z}_{j}$ and $\tilde{\gamma}_{j}$ we denote reconstructed poles and coefficients, respectively, and by the following formulas
@@
min 1.781.78e–$04$ 1.171.17e–$04$ min 4.034.03e–$02$ 2.772.77e–$02$
max 2.112.11e–$03$ 2.462.46e–$03$ max 3.043.04e–$01$ 1.441.44e–$01$
average 1.441.44e–$03$ 8.488.48e–$04$ average 1.111.11e–$01$ 9.559.55e–$02$
In Examples 5.1-5.3, we demonstrate recovery results for three different rational functions and analyze them employing sensitivity analysis as was discussed in Section 4. We utilize $2N$ noisy Fourier coefficients $\hat{r}_{-k}(1+\varepsilon_{k})$, $k=1,\ldots,2N$, to reconstruct poles $z_{n}^{(-)}$, and $2N$ noisy Fourier coefficients $\hat{r}_{k}(1+\varepsilon_{k})$, $k=1,\ldots,2N$, to reconstruct poles $z_{n}^{(+)}$. By $(\varepsilon_{k})_{k=1}^{2N}$ we denote the random noise that is generated in Matlab by $\sigma*$randn(2N), i.e. we consider a multiplicative noise drawn from the standard normal distribution. We set $L=N$ and perform computations for different values of $N$ and $\sigma$. For $\sigma>0$ we perform 10 itterations. To compute sensitivities $\rho_{n}^{(-)}$, $\rho_{n}^{(+)}$ and $\eta_{n}^{(-)}$, $\eta_{n}^{(+)}$, we use formulas (4.14), (4.20) and (4.28), (4.29), respectively. Upper bounds for sensitivities $\rho_{n}^{(-)}$ and $\rho_{n}^{(+)}$ are computed via (4.15) and (4.21), respectively.
Table 3: Reconstruction errors $|z_{j}-\tilde{z}_{j}|$ for poles $z_{j}$ for Example 5.2 for $\sigma=10^{-3}$
@@
min 1.711.71e–$05$ 1.131.13e–$06$ 1.311.31e–$05$ 8.748.74e–$06$
max 2.862.86e–$04$ 2.72.7e–$04$ 1.891.89e–$04$ 1.91.9e–$04$
average 1.441.44e–$04$ 1.191.19e–$04$ 1.041.04e–$04$ 9.439.43e–$05$
Table 4: Reconstruction errors $|z_{j}-\tilde{z}_{j}|$ for poles $z_{j}$ for Example 5.2 for $\sigma=10^{-1}$
@@
min 4.24.2e–$03$ 4.254.25e–$01$
max 1.731.73e–$02$ 1.221.22e–$00$
average 9.79.7e–$03$ 6.356.35e–$01$
Example 5.1.
Let a rational function $r(z)$ as in (1.1) be given by parameters $M=2$ and
$\boldsymbol{z}=(-0.1,-2.1)^{T},\quad\boldsymbol{\gamma}=(0.5,0.5)^{T}$
(see [6, 24]). When $\sigma=0$, i.e. in case of recovery from exact data, we get the following reconstruction errors for $N=2$
[…]
Since we are interested to investigate which pole is the most sensitive to perturbations, we compute reconstruction errors for poles $z_{1}^{(-)}=-0.1$ and $z_{1}^{(+)}=-2.1$ independently and present the results in Table 1. Results for the recovery of the coefficient vector are shown in Table 2.
@@
4 $1.1\mathrm{i}$ 0.035 2.208 3.559
Analyzing numerical results from Table 1, we conclude that for $\sigma=10^{-1}$ our method still reconstructs the locations of the poles quite good. On the other hand, employing more noisy Fourier coefficients also gives us better recovery results. We notice that a pole $z_{1}^{(-)}=-0.1$ is less sensitive to perturbations than a pole $z_{1}^{(+)}=-2.1$. This numerical observation is also confirmed with the sensitivities values. For a pole $z_{1}^{(-)}=-0.1$, we have $\rho_{1}^{(-)}=2\cdot 10^{-1}$ and $\eta_{1}^{(-)}=1.97$, and for a pole $z_{1}^{(+)}=-2.1$, we have $\rho_{1}^{(+)}=9.52\cdot 10^{-1}$ and $\eta_{1}^{(+)}=6.65$. Both unstructured and structured sensitivities are larger for a pole $z_{1}^{(+)}=-2.1$ located outside the unit circle. We compute also vectors (4.27)
$(\boldsymbol{\eta}^{(1)}_{2})^{(-)}=(2.0\cdot 10^{-1},2.0)^{T}\text{ and }(\boldsymbol{\eta}^{(1)}_{2})^{(+)}=(4.2,8.82)^{T},$
which provide us with the additional information that a pole $z_{1}^{(-)}=-0.1$ is the most sensitive to the perturbations of the Fourier coefficient $\hat{r}_{-2}$, and a pole $z_{1}^{(+)}=-2.1$ is the most sensitive to the perturbations of the Fourier coefficient $\hat{r}_{2}$.
Table 7: Reconstruction errors $|z_{j}-\tilde{z}_{j}|$ for poles $z_{j}$ for Example 5.3 for $N=4$
$\sigma=10^{-7}$ $z_{1}^{(-)}=0.2$ $z_{2}^{(-)}=0.5$ $z_{1}^{(+)}=2$ $z_{2}^{(+)}=50$
min 1.581.58e–$08$ 2.92.9e–$08$ 6.076.07e–$09$ 1.751.75e–$03$
max 5.15.1e–$07$ 2.492.49e–$07$ 2.72.7e–$07$ 3.213.21e–$01$
average 2.482.48e–$07$ 1.31.3e–$07$ 1.451.45e–$07$ 1.341.34e–$01$
$\sigma=10^{-5}$ $z_{1}^{(-)}=0.2$ $z_{2}^{(-)}=0.5$ $z_{1}^{(+)}=2$ $z_{2}^{(+)}=50$
min 4.244.24e–$06$ 1.181.18e–$06$ 2.572.57e–$06$ 3.163.16e–$00$
max 6.456.45e–$05$ 2.772.77e–$05$ 2.922.92e–$05$ 9.469.46e+$01$
average 2.312.31e–$05$ 1.131.13e–$05$ 1.51.5e–$05$ 2.022.02e+$01$
@@
We consider the case when poles are located very close to the unit circle. Let a rational function $r(z)$ as in (1.1) of type $(7,8)$, i.e with $M=8$, be given by vectors of poles and coefficients
$\boldsymbol{z}=(0.9,-0.9,0.9\mathrm{i},-0.9\mathrm{i},1.1,-1.1,1.1\mathrm{i},-1.1\mathrm{i})^{T},\quad\boldsymbol{\gamma}=(1,2,3,4,5,6,7,8)^{T},$
respectively. First, we consider recovery from the exact data, i.e for $\sigma=0$. For $N=8$, we obtain the following reconstruction errors
To reconstruct from noisy data, we consider $\sigma=10^{-3}$ and $\sigma=10^{-1}$. The recovery results for each pole $z_{j}$ are presented in Tables 3 and 4, as well as for the coefficient vector in Table 5. Based on these numerical results, we conclude that a pole $z_{1}^{(-)}=0.9$ is the most sensitive to perturbations, which is corroborated by the sensitivities shown in Table 6.
@@
Finally, let a rational function $r(z)$ as in (1.1) of type $(3,4)$, i.e with $M=4$, be given by vectors of poles and coefficients
$\boldsymbol{z}=(0.2,0.5,2,50)^{T},\quad\boldsymbol{\gamma}=(1,1,1,1)^{T},$
Then we consider recovery from noisy data with $\sigma=10^{-7}$ and $\sigma=10^{-5}$, and present the corresponding results in Table 7. Analyzing these numerical results, we can conclude that a pole $z_{2}^{(+)}=50$ is the most sensitive to perturbations. For $\sigma=10^{-5}$ the method does not reconstruct location of this pole correctly. This numerical observation is also predicted by the sensitivities values shown in Table 8. Both unstructured $\rho_{2}^{(+)}=2.204\cdot 10^{3}$ and structured $\eta_{2}^{(+)}=1.577\cdot 10^{4}$ sensitivities are sufficiently larger than sensitivities of other poles. Note, that the locations of the poles $z_{1}^{(-)}=0.2$, $z_{2}^{(-)}=0.5$ and $z_{1}^{(+)}=2$ are reconstructed even for $\sigma=10^{-1}$. We present results regarding recovery of the coefficient vector in Table 9.
@@
min 7.387.38e–$04$ 3.793.79e–$02$
max 7.387.38e–$03$ 6.956.95e–$01$
average 44e–$03$ 4.114.11e–$01$
We compute also vectors (4.27) for poles located inside the unit circle
$(\boldsymbol{\eta}^{(1)}_{4})^{(-)}=(0.555,4.999,13.333,11.111)^{T},\quad(\boldsymbol{\eta}^{(2)}_{4})^{(-)}=(0.222,2.666,9.999,11.111)^{T},$
and outside the unit circle
$(\boldsymbol{\eta}^{(1)}_{4})^{(+)}$ $=(3.472\cdot 10^{-3},0.354,9.375,17.36)^{T},$
$(\boldsymbol{\eta}^{(2)}_{4})^{(+)}$ $=(54.25,2.929\cdot 10^{3},1.106\cdot 10^{4},1.085\cdot 10^{4})^{T}.$
This implies that the poles $z_{1}^{(-)}=0.2$ and $z_{2}^{(-)}=0.5$ are the most sensitive to perturbations of the Fourier coefficients $\hat{r}_{-3}$ and $\hat{r}_{-4}$, respectively, while the poles $z_{1}^{(+)}=2$ and $z_{2}^{(+)}=50$ are the most sensitive to the perturbations of the Fourier coefficients $\hat{r}_{4}$ and $\hat{r}_{3}$, respectively.
Table 10: $\ell_{p}$-norms of $\boldsymbol{\rho}^{(-)}$ and $\boldsymbol{\rho}^{(+)}$ and their upper bounds for Example 5.3 for $p=1,2$
@@
$N=4$ 4.014.01e+$01$ 5.7255.725e+$01$ 2.2092.209e+$03$ 6.8026.802e+$03$
In Subsection 4.2, we derived the upper bounds (4.22)–(4.24) for the $\ell_{2}$- and the $\ell_{1}$-norms of vectors $\boldsymbol{\rho}^{(-)}$ and $\boldsymbol{\rho}^{(+)}$. In order to understand how sharp these estimates are, we compute the exact values for the corresponding norms and their upper bounds as in (4.22)–(4.24) and present the obtained numerical results in Table 10. Comparing these results, we can conclude that the estimates (4.22)–(4.24) are quite sharp.
Table 11: Reconstruction results for Example 5.1
@@
min 4.344.34e–$03$ 1.061.06e–$03$ min 2.372.37e–$02$ 2.272.27e–$02$
max 1.231.23e–$02$ 4.244.24e–$03$ max 2.052.05e+$00$ 7.567.56e–$02$
average 7.647.64e–$03$ 2.712.71e–$03$ average 4.524.52e–$01$ 4.554.55e–$02$
Table 12: Reconstruction results for Example 5.2
@@
min 2.22.2e–$03$ 2.422.42e–$03$ min 1.541.54e–$02$ 9.739.73e–$03$
max 1.371.37e–$02$ 4.864.86e–$02$ max 6.696.69e–$02$ 3.543.54e–$02$
average 6.936.93e–$03$ 1.061.06e–$02$ average 4.424.42e–$02$ 5.755.75e–$02$
In Example 5.4, we discuss the comparison of the performance of our approach with the AAA method for rational approximation. We employ $4N$ noisy function values $r(t_{k})(1+\varepsilon_{k})$ with $t_{k}=\mathrm{e}^{\frac{2\pi\mathrm{i}k}{4N}}$ for $k=0,\ldots,4N-1$, as input information for both methods. By $(\varepsilon_{k})_{k=0}^{4N-1}$ as above we denote the random noise that is generated in Matlab by $\sigma*$randn(4N). The corresponding modification of our method is discussed in Remark 4.1. We refer the interested reader to [15] regarding description of the AAA method. Note only that AAA takes the form of iteration with respect to a type of a rational approximant that is represented at each step in the barycentric form. In case of reconstruction of rational function of type (M1,M)(M-1,M) using exact data, AAA terminates after $M+1$ iteration steps and returns $r$ as
@@
where points $t_{k_{s}}$ are chosen by some greedy algorithm and weights $w_{s}$ are computed by solving a linear least-squares problem over the subset of sample points that have not been selected as support points, $\{t_{0},\ldots,t_{4N-1}\}\setminus\{t_{k_{1}},\ldots,t_{k_{M+1}}\}$ (see also [9, 10, 11] for details). The AAA method has the overall computational cost of $\mathcal{O}(NM^{3})$ operations.
Table 13: Reconstruction results for Example 5.3
@@
We consider again recovery of rational functions $r$ from Examples 5.15.3. In Tables 1113 we present the results for $N=15$, $L=20$ (for the modified method $L\leq 2N$) and different values of $\sigma$.
Analyzing results of Tables 1113, we conclude that for $r$ from Example 5.2 both methods work very well, while for $r$ from Examples 5.1 and 5.3 we get a slightly better recovery applying our method. This is because the Hankel pencil-based approach is more effective for the recovery of poles with certain locations. For Example 5.3 the poles $z_{1}^{(-)}=0.2$, $z_{2}^{(-)}=0.5$ and $z_{1}^{(+)}=2$ are reconstructed equally good by both methods, while the Hankel pencil method shows better performance for the recovery of the pole $z_{2}^{(+)}=50$ than the AAA method (see Table 14).
Table 14: Reconstruction errors $|z_{j}-\tilde{z}_{j}|$ for poles $z_{j}$ from Example 5.3 with the Hankel pencil and the AAA methods